# Temporal Causal Discovery in Dynamic Bayesian Networks Using Federated Learning

### v1

@@ 4.2 Varying Number of Clients
Figure 5: Structure learning for $W$ of a DBN with Gaussian noise for $d=10$ variables, $p=1$, and varying numbers of clients. There are $n=256$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
@@
Figure 6: Structure learning for $W$ of a DBN with Gaussian noise for $d=20$ variables, $p=1$, and varying numbers of clients. There are $n=256$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
@@
From Figures 5 and 6, 2Dbn consistently attains the lowest SHD compared with Best and Ave. For $d=20$, the TPRs of Ave and Best continue to decline as $K$ increases, whereas 2Dbn’s TPR remains higher for all $K\geq 8$. At extremely high distributions, such as $K=64$ with $d=20$, each client has only 4 samples, leading to non-convergence for Ave and Best. Thus, we compare metrics at $K=32$ for a fair assessment. For $d=10$, while Ave achieves the highest TPR among all methods for all $K$, its FDR is much larger than that of the other methods. This suggests that the thresholding strategy used by Ave may lead to a high number of false positives. Nonetheless, 2Dbn’s TPR for $d=10$ remains close to that of Alldata as $K$ grows.
@@
In some cases, when $K$ or $d$ is small, the other baselines may outperform 2Dbn. For example, with $d=10,n=512$, Ave and Best achieve better SHDs than 2Dbn when $K\leq 16$. This is expected since fewer clients and lower dimensionality provide each client with enough samples for accurate independent structure learning. However, as complexity grows (e.g., $d=20$), 2Dbn outperforms the baselines by maintaining a lower FDR and effectively leveraging distributed information. These findings highlight the importance of both method selection and hyperparameter tuning, as well as potential directions for future research in improving the efficiency of information exchange mechanisms for DBN learning.
@@ 5.1 DREAM4
Building on the DREAM4 Challenge [29, 45, 46], our paper focuses on the time-series track of the InSilico_Size100 subchallenge. In this problem, the DREAM4 gene expression data are used to infer gene regulatory networks. The InSilico_Size100 subchallenge dataset contains 5 independent datasets, each consisting of 10 time-series for 100 genes, measured over 21 time steps. We assume that each dataset corresponds to data collected from different hospitals or research centers, thereby motivating our federated approach. More information and data about DREAM4 are available at https://gnw.sourceforge.net/resources/DREAM4%20in%20silico%20challenge.pdf.
@@
Let $X^{g}_{t,r}$ denote the expression level of gene $g$ at time $t\in\{0,1,2,\ldots,20\}$ in replicate $r\in\{0,1,2,\ldots,R\}$. Note that $R$ depends on the dataset used — for instance, if only one dataset is used, then $R=10$. Consequently, $X_{t,r}\in\mathbb{R}^{100}$ and $X_{t}\in\mathbb{R}^{R\times 100}$. In this experiment, we set $p=1$, aligning with the VAR method proposed in the DREAM4 Challenge by Lu et al. [26]. The federated setting of our 2Dbn approach includes $K=5$ clients, each with $R=2$ replicates. Thus, each client contains a time-series dataset for 100 genes over 42 time steps. We found that small regularization parameters, $\lambda_{w}=\lambda_{a}=0.0025$, work well for all DREAM4 datasets. Lu et al. [26] evaluated various methods for learning these networks, including approaches based on Mutual Information(MI), Granger causality, dynamical systems, Decision Trees, Gaussian Processes (GPs), and Dynamic Bayesian Networks. Notably, we did not compare DYNOTEARS in this study, as its source code is not publicly available, and our implemented version did not achieve optimal results. The comparisons were made using AUPR (Area Under the Precision-Recall Curve) and AUROC (Area Under the Receiver Operating Characteristic Curve) across the 5 datasets. From Table. 1, the GP-based method outperforms all others, achieving the highest mean AUPR (0.208) and AUROC (0.728). In terms of AUPR, our 2Dbn approach outperforms the TSNI (ODE-based) method in its non-federated version. Furthermore, the AUPR of 2Dbn is comparable to those of Ebdnet (DBN-based), GCCA (VAR-based), and ARACNE (MI-based) approaches. Importantly, the mean AUROC of 2Dbn surpasses those of GCCA (VAR-based), ARACNE (MI-based), and TSNI (ODE-based) methods, and is close to those of Ebdnet and VBSSMa (DBN-based).
@@
Overall, 2Dbn is comparable to other non-federated benchmarks. However, the GP-based method still achieves superior results in both AUPR and AUROC. One possible reason for this superiority is that the GP-based approach may more effectively capture nonlinear relationships and complex temporal dynamics, which can be challenging in a federated setting where data distributions and conditions vary across sources. Unsurprisingly, the GP-based method outperforms all approaches because Gaussian Processes naturally handle nonlinear dynamics and are more adaptive to varying, complex relationships. In contrast, the 2Dbn model, although suitable for distributed data, might impose stronger parametric assumptions or face computational constraints that limit its ability to capture subtle nonlinear interactions. Thus, we believe GP-based methods could be extended to a distributed approach for structure learning, making them more applicable to real-world problems. We discuss this further in Section 6.
@@ 5.2 Functional Magnetic Resonance Imaging (FMRI)
For the federated setting, we partition the 200 time steps among 5 clients ($K=5$). Detailed information and descriptions of the data are available at https://www.fmrib.ox.ac.uk/datasets/netsim/index.html. In our experiments, we evaluate the proposed method for $d\in\{5,10,15\}$. For $d=5$, we use the 3rd, 6th, 9th, 12th, and 15th subjects from Sim-1.mat. For $d=10$, we use the 2nd, 4th, 6th, 8th, and 10th subjects from Sim-2.mat. Finally, for $d=15$, we use the 1st, 3rd, 5th, 7th, and 9th subjects from Sim-3.mat. Further details of this experiment are provided in the supplementary materials (see §7.5). We compare our method to the economy Statistical Recurrent Units (eSRU) proposed by Khanna and Tan [19] for inferring Granger causality, as well as existing methods based on a Multilayer Perceptron (MLP), a Long Short-Term Memory (LSTM) network [48], and a Convolutional Neural Network (CNN)-based model, the Temporal Causal Discovery Framework (TCDF) [34], on multivariate time series data for $d=15$, as examined by Khanna and Tan [19]. As shown in Tab.2, our proposed 2DBNs achieve an AUROC of 0.74, outperforming the LSTM-based approach and approaching the performance of the CNN-based TCDF method. Even though 2DBNs do not surpass the MLP-based and eSRU-based methods, it is notable that the Federated version of our approach outperforms or closely matches several established non-Federated benchmarks. This outcome is reasonable, as the MLP and eSRU methods rely on deep architectures adept at modeling complex structural dependencies. Notably, our 2DBN method provides a new perspective on this problem by ensuring data security through its Federated approach. This capability is particularly important in scenarios involving sensitive or distributed datasets, as it allows for effective analysis without compromising privacy or data integrity. We have further elaborated on this aspect in Sec.6.

### v2

@@ 4.1.2 Varying Number of Clients
Figure 6: Structure learning for $W$ of a DBN with Gaussian noise for $d=10$ variables, $p=1$, and varying numbers of clients. There are $n=256$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$. Each metric value indicates the mean performance across 10 different simulated datasets
@@
Figure 7: Structure learning for $W$ of a DBN with Gaussian noise for $d=20$ variables, $p=1$, and varying numbers of clients. There are $n=256$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$. Each metric value indicates the mean performance across 10 different simulated datasets
@@
From Figures 6 and 7, FDBNL consistently attains the lowest SHD compared with Best and Ave. For $d=20$, the TPRs of Ave and Best continue to decline as $K$ increases, whereas FDBNL’s TPR remains higher for all $K\geq 8$. At extremely high distributions, such as $K=64$ with $d=20$, each client has only 4 samples, leading to non-convergence for Ave and Best. Thus, we compare metrics at $K=32$ for a fair assessment. For $d=10$, while Ave achieves the highest TPR among all methods for all $K$, its FDR is much larger than that of the other methods. This suggests that the thresholding strategy used by Ave may lead to a high number of false positives. Nonetheless, FDBNL’s TPR for $d=10$ remains close to that of Alldata as $K$ grows.
@@
In some cases, when $K$ or $d$ is small, the other baselines may outperform FDBNL. For example, with $d=10,n=512$, Ave and Best achieve better SHDs than FDBNL when $K\leq 16$. This is expected since fewer clients and lower dimensionality provide each client with enough samples for accurate independent structure learning. However, as complexity grows (e.g., $d=20$), FDBNL outperforms the baselines by maintaining a lower FDR and effectively leveraging distributed information. These findings highlight the importance of both method selection and hyperparameter tuning, as well as potential directions for future research in improving the efficiency of information exchange mechanisms for DBN learning.
4.2 Personalized Federated Dynamic Bayesian Network Learning
In this section, we first compare PFDBNL to FDBNL. For this example, we generate six pairs of $(W,A)$ for each client. Then, we simulate 10 datasets with $n_{k}=30$ for the $k$-th client. We compute the mean mSHD, mTPR, and mFDR after applying PFDBNL and FDBNL to each of these datasets.
From Table 1, We observe that the mSHD of $W$ and $A$ in PFDBNL is half that of $W$ and $A$ in FDBNL. Furthermore, the mTPR of PFDBNL is twice as high as that of FDBNL. For mFDR, we observe that PFDBNL’s value is slightly lower than FDBNL’s, but the mFDR decreases by a factor of two. Hence, PFDBNL significantly improves performance across all evaluation metrics compared to FDBNL. To the best of our knowledge, no existing method is capable of learning Dynamic Bayesian Networks (DBNs) from heterogeneous time-series data. Given this gap, PFDBNL represents the first attempt to address this challenge and demonstrates its effectiveness in improving accuracy.
Metric PFDBNL FDBNL
mSHD $W$ 6.2 10.2
$A$ 5.4 12.7
mTPR $W$ 0.64 0.35
$A$ 0.51 0.24
mFDR $W$ 0.55 0.60
$A$ 0.19 0.33
Table 1: Comparison between FDBNL and PFDBNL with $d=5$, $K=6$, $n_{k}=30$ for $k\in K$. The expected degree of a node is 4 (both direction) with $p=1$. Each metric value indicates the mean performance across 10 different simulated datasets
Data Generation & Settings.
For PFDBNL, we generate data following the Structural Equation Model described in Eq. (2). This involves three steps: (1) Constructing weighted graphs $G_{W_{k}}$ and $G_{A_{k_{i}}}$ for each client. (2) Creating data matrices $X_{\text{client}}$ and $Y_{\text{client}}$ aligned with these graphs for each client. (3) Applying PFDBNL and evaluating its performance. Steps (1) and (2) follow the same data generation procedure as FDBNL, but partitioning of the dataset is not required. We use Gaussian noise with a standard deviation of 1 and set the base of the exponential decay of inter-slice weights to $\eta=1.5$. For the hyperparameters $\lambda_{W}$, $\lambda_{A}$, and $\mu$, we found that setting $\lambda_{W}=\lambda_{A}=0.1$ works well for $d=\{5,10\}$, while $\lambda_{W}=\lambda_{A}=0.01$ is effective for $d=\{15,20\}$ with $\mu=0.1$. We set $\phi_{1}=1.6$ and $\phi_{2}=1.1$, with initial values $\rho_{1}=\rho_{2}=1$. The initial Lagrange multipliers are set to zero. The data shape is the same as in the FDBNL experiment, but each client has a different $W_{k},A_{k_{i}}$.
Experiment Settings. We consider the same two types of experimental settings as in the FDBNL experiment. First, we fix the number of clients $K$ and increase the number of variables $d$ while maintaining a consistent sample size $n$. This allows us to assess how well the methods scale with increasing dimensionality. Second, we fix the total sample size at $n=512$ and vary the number of clients $K$ from 2 to 32. This setting evaluates the adaptability and robustness of the methods as the data becomes more distributed.
4.2.1 Varying Number of Variables
In this section, we focus on learning of $W_{k}$ and $A_{k_{i}}$ with each $n_{k}=Kd$ samples for $d\in\{5,10,15,20\}$ across $K=6$ clients. Similarly to the example, we firstly generate $W_{k},A_{k_{1}}$ by using $p=1$ for all clients. Then we generate the dataset with $n_{k}$ sample sizes for $k$-th client. Furthermore, we exam our PFDBNL under two different connectivity level (one for low connectivity and one for high connectivity). We do not focused on mixed case in this study and leave it for future research direction as discussed in §6. For low connectivity, we set the expectation of degree of a node below half of number of variable as $\lfloor\frac{d}{2}\rfloor-\frac{d}{5}$. For high connective, we set expectation of degree of a node as $d-\frac{d}{5}$.
Metric 5 10 15 20
mSHD $W$ 4.4 16.8 44.7 79.2
$A$ 2.8 10.9 29.6 32.2
mTPR $W$ 0.61 0.49 0.55 0.53
$A$ 0.81 0.53 0.31 0.31
mFDR $W$ 0.72 0.67 0.51 0.53
$A$ 0.35 0.47 0.23 0.34
Table 2: Comparison of metrics (mSHD, mTPR, mFDR) across dimensions $d=5,10,15,20$ for low connectivity. Each metric value indicates the mean performance across 10 different simulated datasets
Metric 5 10 15 20
mSHD $W$ 6.2 13.6 52.0 110.0
$A$ 5.4 8.3 41.0 31.0
mTPR $W$ 0.64 0.64 0.71 0.76
$A$ 0.51 0.52 0.31 0.27
mFDR $W$ 0.55 0.54 0.44 0.59
$A$ 0.19 0.21 0.36 0.19
Table 3: Comparison of metrics (mSHD, mTPR, mFDR) across dimensions $d=5,10,15,20$ for high connectivity. Each metric value indicates the mean performance across 10 different simulated datasets
From Tables 2 and 3, we observe that the mSHD for the low-connectivity case is much lower than in the high-connectivity case for $W$. However, the mSHD remains similar across all cases for $A$. For mTPR, we observe that performance in the high-connectivity scenario is significantly better than in the low-connectivity scenario. More importantly, while the mTPR for high-connectivity does not decrease with an increasing number of variables for $W$, it does decrease for $A$. Additionally, the mTPR for $A$ remains similar for $d\in\{10,15,20\}$, but in the low-connectivity case, it is 0.81, which is significantly higher than 0.51 in the high-connectivity case. This is not surprising, as identifying relationships in sparse networks is more challenging. For mFDR of $A$, we observe that it remains around 0.2 for $d\in\{5,10,20\}$ in the high-connectivity case. However, for the low-connectivity case, the mFDR increases to approximately 0.3. For $W$, the mFDR remains stable at around 0.54 in the high-connectivity scenario, which is significantly lower than in the low-connectivity case when $d\in\{5,10\}$. Overall, the mFDR in the high-connectivity case shows better results.
We further explore the partial participation case under $d=10$. Since we only have six clients, we set the number of participating clients to $j=1,2,3,4,5$, while keeping all other settings unchanged. As shown in Table 4, the results remain stable regardless of the number of participating clients. The mSHD is approximately 13 to 14 for $W$ and around 8 for $A$. The mTPR remains about 0.60 for $W$ and 0.56 for $A$. The mFDR follows a similar pattern.
Metric $j=1$ $j=2$ $j=3$ $j=4$ $j=5$
mSHD $W$ 14.2 13.6 14.3 14.9 14.5
$A$ 8.1 8.3 8.7 7.7 8.5
mTPR $W$ 0.59 0.61 0.60 0.58 0.60
$A$ 0.56 0.55 0.55 0.58 0.56
mFDR $W$ 0.50 0.49 0.53 0.55 0.54
$A$ 0.22 0.21 0.23 0.23 0.23
Table 4: Partial participating case for $d=10$. Each metric value indicates the mean performance across 10 different simulated datasets
4.2.2 Varying Number of Clients
We now examine scenarios where a fixed total number of samples is distributed among varying numbers of clients. For $d\in\{10,20\}$, we generate $n=512$ samples in total. These samples are evenly distributed across $K\in\{2,4,8,16,32\}$ clients, meaning that each client receives $n_{k}=\{256,128,64,32,16\}$ samples, respectively. The mean degree of the intra-slice DAG is 4, counting edges in both directions. This may be a challenging setting because we need to learn different DBNs for 32 clients with only 16 samples per client.
Metric $K=2$ $K=4$ $K=8$ $K=16$ $K=32$
mSHD $W$ 11.3 20.6 17.4 18.2 19.4
$A$ 6.7 10.5 10.7 10.8 14.8
mTPR $W$ 0.68 0.44 0.51 0.51 0.49
$A$ 0.63 0.52 0.48 0.48 0.36
mFDR $W$ 0.49 0.62 0.54 0.54 0.59
$A$ 0.17 0.31 0.23 0.28 0.37
Table 5: Comparison of metrics (mSHD, mTPR, mFDR) across dimensions $K=2,4,8,16,32$ for $d=10$. Each metric value indicates the mean performance across 10 different simulated datasets
Metric $K=2$ $K=4$ $K=8$ $K=16$ $K=32$
mSHD $W$ 56.2 57.5 54.3 52.1 56.1
$A$ 33.2 36.1 40.8 55.6 56.3
mTPR $W$ 0.45 0.46 0.46 0.34 0.32
$A$ 0.38 0.31 0.33 0.25 0.17
mFDR $W$ 0.40 0.45 0.37 0.49 0.51
$A$ 0.18 0.16 0.16 0.21 0.27
Table 6: Comparison of metrics (mSHD, mTPR, mFDR) across dimensions $K=2,4,8,16,32$ for $d=20$. Each metric value indicates the mean performance across 10 different simulated datasets
From Tables 5 and 6, we observe that PFDBNL consistently achieves mSHD values around 10–20 for $W$ and approximately 10 for $A$ when $d=10$. Notably, the mTPR of $W$ remains relatively stable as $K$ increases. For instance, at $K=32$, $W$ and $A$ obtain mTPRs of 0.49 and 0.36, respectively. This result is encouraging, even in scenarios where the number of clients ($K$) exceeds the number of samples per client. Meanwhile, the mFDR remains around 0.2–0.3 for $A$ but roughly 0.5 for $W$, suggesting that $A$ is more conservative, with fewer false discoveries.
For $d=20$, although $W$ stabilizes at an mSHD of around 55, the performance of $A$ deteriorates from an mSHD of approximately 30 to 50 as $K$ increases, indicating greater sensitivity to high-dimensional settings. A similar pattern is observed in the mTPR: $W$ achieves approximately 0.45 when $K\in\{2,4,8\}$ but declines to about 0.3 for larger $K$. This drop is expected, as the sample size per client decreases while the number of parameters in the ground-truth matrix remains large. Overall, however, PFDBNL maintains stable and reasonable performance across different values of $K$, including the challenging scenario of $d=20,K=32,n_{k}=16$.
@@ 5.1 DREAM4
Building on the DREAM4 Challenge [33, 53, 54], our paper focuses on the time-series track of the InSilico_Size100 subchallenge. In this problem, the DREAM4 gene expression data are used to infer gene regulatory networks. The InSilico_Size100 subchallenge dataset contains 5 independent datasets, each consisting of 10 time-series for 100 genes, measured over 21 time steps. We assume that each dataset corresponds to data collected from different hospitals or research centers, thereby motivating our federated approach. More information and data about DREAM4 are available at https://gnw.sourceforge.net/resources/DREAM4%20in%20silico%20challenge.pdf
@@
Let $X^{g}_{t,r}$ denote the expression level of gene $g$ at time $t\in\{0,1,2,\ldots,20\}$ in replicate $r\in\{0,1,2,\ldots,R\}$. Note that $R$ depends on the dataset used—for instance, if only one dataset is used, then $R=10$. Consequently, $X_{t,r}\in\mathbb{R}^{100}$ and $X_{t}\in\mathbb{R}^{R\times 100}$. In this experiment, we set $p=1$, aligning with the VAR method proposed in the DREAM4 Challenge by Lu et al. [30]. The federated setting of our FDBNL approach includes $K=5$ clients, each with $R=2$ replicates. Thus, each client contains a time-series dataset for 100 genes over 42 time steps. We found that small regularization parameters, $\lambda_{W}=\lambda_{A}=0.0025$, work well for all DREAM4 datasets. For our PFDBNL approach, we assume each dataset to be a client. Consequently, each client has 200 time steps with $d=100$ and $K=5$. We used regularization parameters $\lambda_{W}=\lambda_{A}=0.001$ and $\mu=0.1$.
Lu et al. [30] evaluated various methods for learning these networks, including approaches based on Mutual Information (MI), Granger causality, dynamical systems, Decision Trees, Gaussian Processes (GPs), and Dynamic Bayesian Networks (DBNs). Notably, we did not compare DYNOTEARS in this study, as its source code is not publicly available, and our implemented version did not achieve optimal results. The comparisons were made using AUPR (Area Under the Precision-Recall Curve) and AUROC (Area Under the Receiver Operating Characteristic Curve) across the five datasets.
From Table 7, the GP-based method outperforms all others, achieving the highest mean AUPR (0.208) and AUROC (0.728). In terms of AUPR, our FDBNL approach outperforms the TSNI (ODE-based) method in its non-federated version. Furthermore, the AUPR of FDBNL is comparable to those of Ebdnet (DBN-based), GCCA (VAR-based), and ARACNE (MI-based) approaches. Importantly, the mean AUROC of FDBNL surpasses those of GCCA (VAR-based), ARACNE (MI-based), and TSNI (ODE-based) methods, and is close to those of Ebdnet and VBSSMa (DBN-based). We also directly applied FDBNL, marked as FDBNL-d, to all datasets using the same settings as PFDBNL. From the table, we can see that if we directly use FDBNL, the results are not satisfactory. Therefore, we partitioned each dataset to ensure i.i.d. conditions for the use of FDBNL. We then tested PFDBNL, which performs well in heterogeneous situations. For PFDBNL, the AUPR is 0.034, which outperforms the TSNI (ODE-based) method in its non-personalized version. Additionally, the AUPR of FDBNL is only slightly greater than that of PFDBNL. Moreover, PFDBNL is also comparable to Ebdnet (DBN-based), GCCA (VAR-based), and ARACNE (MI-based) approaches. More importantly, the mean AUROC of PFDBNL surpasses that of ARACNE (MI-based) and is close to those of TSNI (ODE-based) and GCCA (VAR-based) methods.
@@
Overall, FDBNL and PFDBNL are comparable or even sometimes superior to other non-federated benchmarks. However, the GP-based method still achieves superior results in both AUPR and AUROC. One possible reason for this superiority is that the GP-based approach may more effectively capture nonlinear relationships and complex temporal dynamics, which can be challenging in a federated setting where data distributions and conditions vary across sources. Unsurprisingly, the GP-based method outperforms all approaches because Gaussian Processes naturally handle nonlinear dynamics and are more adaptive to varying, complex relationships. In contrast, the FDBNL and PFDBNL models, although well-suited for distributed data, may impose stronger parametric assumptions or face computational constraints that limit their ability to capture subtle nonlinear interactions. Thus, we believe GP-based methods could be extended to a distributed approach for structure learning, making them more applicable to real-world problems. We discuss this further in Section 6.
@@ 5.2 Functional Magnetic Resonance Imaging (FMRI)
For the FDBNL setting, we partition the 200 time steps among five clients ($K=5$). Detailed information and descriptions of the data are available at https://www.fmrib.ox.ac.uk/datasets/netsim/index.html. In our experiments, we evaluate the proposed method for $d\in\{5,10,15\}$. For $d=5$, we use the 3rd, 6th, 9th, 12th, and 15th subjects from Sim-1.mat. For $d=10$, we use the 2nd, 4th, 6th, 8th, and 10th subjects from Sim-2.mat. For $d=15$, we use the 1st, 3rd, 5th, 7th, and 9th subjects from Sim-3.mat. For the PFDBNL setting, we do not apply any partitioning; instead, we consider each subject as a client and we aaply our PFDBNL to 5 subjects. Thus, for $d\in\{5,10,15\}$, we have five clients, each containing 200 time steps. For $d=5$, we use the 2nd, 4th, 6th, 8th, and 10th subjects from Sim-1.mat. For $d=\{10,15\}$, we use the 1st, 3rd, 5th, 7th, and 9th subjects from Sim-2.mat and Sim-3.mat. To avoid any confusion, the difference between the two methods is as follows: (1) FDBNL is applied to each subject individually, where each subject is partitioned into five clients. As a result, we need to run five repetitions. (2) PFDBNL is directly applied to the five subjects without partitioning. Further details of this experiment are provided in the supplementary materials (see §7.6).
We compare our method to the Economy Statistical Recurrent Units (eSRU) model proposed by Khanna and Tan [24] for inferring Granger causality, as well as existing methods based on a Multilayer Perceptron (MLP), a Long Short-Term Memory (LSTM) network [56], and a Convolutional Neural Network (CNN)-based model, the Temporal Causal Discovery Framework (TCDF) [39], on multivariate time series data for $d=15$, as examined by Khanna and Tan [24]. As shown in Table 8, our proposed FDBNL achieves an AUROC of 0.74, outperforming the LSTM-based approach and approaching the performance of the CNN-based TCDF method. Although FDBNL does not surpass the MLP-based and eSRU-based methods, it is noteworthy that its federated version outperforms or closely matches several established non-federated benchmarks. Similarly, our proposed PFDBNL also achieves an AUROC of 0.74, demonstrating superior performance over the LSTM-based approach and approaching the CNN-based TCDF method. Notably, PFDBNL can be directly applied to multiple subjects, significantly reducing the computational cost associated with data preparation. While PFDBNL does not exceed the MLP-based and eSRU-based methods, its personalized version consistently outperforms or closely aligns with several established non-personalized benchmarks. This outcome is expected, as the MLP and eSRU methods rely on deep architectures that excel at modeling complex structural dependencies.
Importantly, our FDBNL and PFDBNL methods offer a new perspective on this problem by ensuring data security through their federated and personalized approach. This capability is particularly crucial in scenarios involving sensitive or distributed datasets, as it enables effective analysis without compromising privacy or data integrity. We further elaborate on this aspect in Section 6.